solve for x
y = (-2/3x) + 8

Answer by richwmiller(17219)

y=-(2/3)*x+8
y-8=-(2/3)*x
(y-8)/-(2/3)=x
-(y-8)*3/2=x
-(3/2)(y-8)=x